OpenStudy (gecko):

Integrate from 0 to pi/8: (sec^2(2x))/(tan(2x)+1) Thanks so much in advance!

OpenStudy (anonymous):

put \[u=\tan(2x)+1\] \[du=2\sec^2(2x)dx\]

OpenStudy (gecko):

O: Ahhh I think I've got it lol. Wasn't sure if it was u-sub since we just started with it in class.. Thanks satellite! I believe this isn't the first time you've saved me (:

OpenStudy (anonymous):

\[u(0)=\tan(0)+1=1\] \[u(\frac{\pi}{8})=\tan(\frac{\pi}{4})+1=2\]

OpenStudy (anonymous):

yw. i think the result will be \[\int_1^2\frac{1}{2}\frac{1}{u}du=\frac{1}{2}\ln(2)\]

OpenStudy (anonymous):

or even \[\ln(\sqrt{2})\] if you prefer

OpenStudy (gecko):

Thank you much (: Would it be -(1/2)ln(2) since you subtract with the upper bound?

OpenStudy (anonymous):

oh no. it is F(b)-F(a)

OpenStudy (anonymous):

upper minus lower yes?

OpenStudy (gecko):

Ohh derp :P You're right. I wrote it down wrong lol. Thanks!

OpenStudy (anonymous):

yw